# A Generalized Difference-in-Differences Estimator for Unbiased Estimation of Desired Estimands from Staggered Adoption and Stepped-Wedge Settings

### v1

@@ 2.6 Minimum variance estimator
For settings where there are many unbiased estimators of this form, the investigator would like to select the one with the lowest variance. Again, treating the weights as fixed (i.e., independent of the outcomes), the variance of the estimator for any $\bm{w}$ is given by:
$Var(\hat{\theta}|\bm{w})=Var(\bm{w}^{T}\bm{d})=Var(\bm{w}^{T}\bm{A}\bm{y})=\bm{w}^{T}\bm{A}Var(\bm{y})\bm{A}^{T}\bm{w}.$ (6)
The conversion from the vector of two-by-two DID estimators $\bm{d}$ to the outcome vector $\bm{y}$ is used here since the structure of the correlation among different two-by-two DID estimators is not trivial. In general, $Var(\bm{y})$ will not be known a priori. For the design and selection of the weights, then, a working covariance matrix can be used, denoted by $\bm{M}$. The “working” variance of the estimator is then estimated for a given $\bm{M}$ by:
@@
Denote by $\bm{w}^{*}$ the weight vector that minimizes $Var(\hat{\theta}|\bm{w},\bm{M})$ under the constraint $\bm{F}^{T}\bm{w}^{*}=\bm{v}$.
Note that the variance can then be estimated in the design phase a priori using $\bm{M}$. Since $\bm{w}^{*}$ is selected among weights that give unbiased estimation regardless of $\bm{M}$, misspecification of $\bm{M}$ results only in reduced efficiency, not bias. Moreover, even for variance minimization, the working covariance structure $\bm{M}$ does not need to be specified exactly, only up to a constant factor. The features of $\bm{M}$ that are necessary for the selection of the weights are the correlation within (and, if applicable, between) units across periods and the relative variances of the units and, if applicable, time periods. The working variance matrix can be written as follows:
@@
where $\bm{M}_{r}$ is the working correlation matrix and $\bm{m}_{v}$ is the vector of (relative) variances of the individual observations. If the units are independent, then $\bm{M}_{r}$—and thus $\bm{M}$—will be block-diagonal, with non-zero entries only for the within-unit covariances only. Several common structures could then be specified for the within-unit covariances:
  • Independence: assuming independence of observations would require $\bm{M}_{r}$ to be the identity matrix. This would generate a diagonal $\bm{M}$, where the diagonal entries indicate the relative variances of the different time periods and units.
  • Exchangeable/compound symmetric: assuming compound symmetry in the within-unit correlation would require $\bm{M}_{r}$ to be a block-diagonal matrix, with the non-zero off-diagonal entries all equal to the intra-unit (intra-cluster) correlation coefficient, $\rho$.
@@
If the units moreover are exchangeable (e.g., clusters prior to randomization in a SWT), then all of the blocks will be equivalent to one another. More complex variance structures can be implied by specific outcome models as well (see, e.g., Hooper et al. (2016); Kasza, Hemming, Hooper, Matthews & Forbes (2019)).
@@ 2.7 Estimation and inference
The estimator is then given by:
$\hat{\theta}^{*}=\bm{w}^{*T}\bm{y}.$ (9)
@@
$Var(\hat{\theta}^{*})=\bm{w}^{*T}\bm{A}\widehat{Var}(\bm{y})\bm{A}^{T}\bm{w}^{*}.$ (10)
Alternatively, inference can proceed using permutation inference (akin to randomization tests in the SWT case as in Wang & De Gruttola (2017), Thompson et al. (2018), and Kennedy‐Shaffer et al. (2020), or placebo tests in the staggered adoption case as in Hagemann (2019), MacKinnon & Webb (2020), Shaikh & Toulis (2021), and Roth & Sant’Anna (2023)).
@@ 2.8 Algorithm for construction of estimator
In summary, the estimator $\hat{\theta}$ can be constructed for any setting by the following process.
@@
    Based on the number of periods $J$ and the number of units $N$, identify the matrix $\bm{A}$ that converts the observations $\bm{y}$ to the two-by-two DID estimators $\bm{d}$. Note that for consistency, the units should be numbered from the earliest treatment adopter to the latest.
@@
    Find the space of vectors $\bm{w}$ such that $\bm{F}^{T}\bm{w}=\bm{v}$ and, if relevant, restrict to the set of weights that give unique estimators.
@@
    Determine an appropriate working covariance matrix $\bm{M}$ using subject-matter knowledge, pilot study data, or an assumed outcome model. Note that this only needs to be specified up to a constant scalar factor.
@@ 3 Toy Example
I consider several different assumption settings.
3.1 Assumption (5): Homogeneity
@@
  • If $x+y=1$, then the estimator is only the type-2 estimator $D_{1,2,1,2}$, the “clean” comparison as described by Goodman-Bacon (2021) or the “crossover” estimator CO-1 proposed by Kennedy‐Shaffer et al. (2020);
@@
Note that the purely vertical estimator described by Matthews & Forbes (2017) and Thompson et al. (2018) cannot be constructed as a weighted average of the two-by-two DID estimators.
If we assume independent and homoskedastic observations, then we can use $\bm{M}=\bm{I}_{6}$, the six-by-six identity matrix. The constrained minimization of the variance, then, gives the solution $x+y=1/2$ and the estimator $\hat{\theta}^{*}=\frac{D_{1,2,1,2}-D_{1,2,2,3}}{2}$. Note that this corresponds to the “centrosymmetrized” estimator from Bowden et al. (2021) which, as proven therein, has the minimum variance under these assumptions. The same estimator minimizes the variance under a working correlation assumption that is compound symmetric or AR(1), regardless of the correlation value, as long as it is exchangeable and independent across units.
3.2 Assumption (4): Calendar-Time Heterogeneity
@@
3.3 Assumption (3): Exposure-Time Heterogeneity
@@ 4 Data Example: Stepped Wedge Trial
This study randomized fourteen clusters into seven treatment sequences, with observations in eight months. The first sequence, consisting of two clusters, began the intervention in the second month; each subsequent month, two more clusters began the intervention (Durovni et al. (2014)). The patient outcomes—unsuccessful treatment of the patient (i.e., any outcome other than “cure and treatment completion without evidence of failure” (Trajman et al. (2015))—were retrieved using the Brazilian national tuberculosis information system over one year after diagnosis.
The original analysis found an insignificant reduction in this composite outcome: an odds ratio of 0.93 with a 95% confidence interval of 0.79–1.08 (Trajman et al. (2015)). Accounting for time trends in the analyses, however, re-analysis by Thompson et al. (2018) using the purely vertical non-parametric within-period method found a significant reduction in these events: an odds ratio of 0.78 with a p-value of 0.02. Similarly significant results were found by Kennedy‐Shaffer et al. (2020) using the crossover method, which estimated an odds ratio of 0.703 with a permutation-based p-value of 0.046. These results indicate the possible impact that different methods of accounting for time trends and treatment heterogeneity can have.
I re-analyze these data using several versions of the generalized DID estimator presented here. I first consider four estimators providing summary treatment effects:
@@
  • $\sum_{j=2}^{7}\hat{\theta}^{(4)}_{j}$, the average period-specific treatment effect across months 2–7 of the trial under calendar-time heterogeneity (assumption S4);
  • $\sum_{a=1}^{7}\hat{\theta}^{(3)}_{a}$, the average time-on-treatment effect across exposure times 1–7 under exposure-time heterogeneity (assumption S3); and
  • $\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{j,a}$, the average across all identifiable calendar- and exposure-time combinations of the trial under calendar- and exposure-time heterogeneity (assumption S2).
Note that the period-specific treatment effects in month 8 (the last month of the study) are not identifiable since all clusters were in the treated condition at that point. These clusters do contribute to estimation of $\hat{\theta}^{(5)}$ and $\hat{\theta}^{(3)}_{7}$, however, under assumptions S5 and S3, as those assumptions allow us to borrow information across different calendar times in estimating common treatment effects. The odds ratio estimates and permutation-test p-values (using 1000 permutations) are shown in Table 3. All results shown were calculated using an exchangeable (within-cluster) variance structure with intracluster correlation coefficient $\rho=0.003$ as estimated by Thompson et al. (2018). Results with independent and auto-regressive variance structures are quite similar and shown in Appendix C.
@@
S5 $\hat{\theta}^{(5)}$ 0.771 0.009 -7.35% 0.015
S4 $\sum_{j=2}^{7}\hat{\theta}^{(4)}_{j}$ 0.784 0.031 -6.81% 0.052
S3 $\sum_{a=1}^{7}\hat{\theta}^{(3)}_{a}$ 0.828 0.397 -5.57% 0.423
S2 $\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{j,a}$ 0.801 0.160 -6.11% 0.243
@@
To understand the form of the estimator, we can also examine the weights it gives to each observation (note: the weights to each two-by-two DID estimator are also available, but less interpretable as many different weight vectors can yield the same weights on the observations). The weights for the estimator $\hat{\theta}^{(5)}$, again using an exchangeable variance structure, are shown in the heat map in Figure 1. Note that the clusters are ordered by the sequence of treatment adoption, and pairs of clusters (e.g., 1 and 2) have nearly identical weights because they are in the same sequence and assumed to be exchangeable. These weights can be compared to the information content of specific cells under different heterogeneity and correlation assumptions as proposed by Kasza & Forbes (2019) and Kasza, Taljaard & Forbes (2019). Note heurestically that in this case, the largest (in magnitude) weights are near the diagonal, followed by the top-right and bottom-left cells, which are also the highest information-content cells as shown in Figure 4 of Kasza & Forbes (2019).
@@
This method provides flexibility to investigate more specific treatment effects in the heterogeneous settings as well. Considering calendar-time effects under assumption S4, for example, one can estimate the treatment effect in each of months 2 through 7, $\hat{\theta}^{(4)}_{2},\ldots,\hat{\theta}^{(4)}_{7}$. For additional robustness against exposure-time heterogeneity, one can also estimate month-specific treatment effects under assumption S2 by averaging across exposure times within each of those months: $\hat{\theta}^{(2)}_{j}\equiv\frac{1}{j-1}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{a,j}$ for all $j=2,\ldots,7$. The estimation and permutation test results on both the multiplicative (odds ratio) and additive (risk difference) scales are shown in Table 4. These results can be compared to the results in Table 2 of Thompson et al. (2018), which reports period-specific risk differences estimated using the purely vertical non-parametric within-period estimator (note that those are labelled by treatment periods, one less than the corresponding month numbers given here). Differences between the methods arise as the results presented here use horizontal information as well, but target the same estimand. Note that no p-values are adjusted for multiple testing.
Table 4: Odds ratio and risk difference estimates and permutation test p-values for generalized DID estimators of calendar-time treatment effects of XpertMTB/RIF testing on tuberculosis outcomes in Brazil, 2012.

### v3

@@ 2.6 Minimum variance estimator
For settings where there are many unbiased estimators of this form, I propose using the one with the lowest estimated variance. Again, treating the weights as fixed (i.e., independent of the outcomes), the variance of the estimator for any $\bm{w}$ is given by:
$Var(\hat{\theta}|\bm{w})=Var(\bm{w}^{T}\bm{d})=Var(\bm{w}^{T}\bm{A}\bm{y})=\bm{w}^{T}\bm{A}Var(\bm{y})\bm{A}^{T}\bm{w},$ (7)
written in terms of $Var(\bm{y})$ rather than $Var(\bm{d})$ because the correlation among different two-by-two DID estimators is not trivial. In general, $Var(\bm{y})$ will not be known a priori. For the design and selection of the weights, then, a working covariance matrix can be used, denoted by $\bm{M}$. The working variance of the estimator is then estimated for a given $\bm{M}$ by:
@@
Let $\bm{w}^{*}$ be the weight vector that minimizes $Var(\hat{\theta}|\bm{w},\bm{M})$ under the constraint $\bm{F}^{T}\bm{w}^{*}=\bm{v}$.
The variance can then be estimated in the design phase a priori using $\bm{M}$. Since $\bm{w}^{*}$ is selected among weights that give unbiased estimation regardless of $\bm{M}$, misspecification of $\bm{M}$ results only in reduced efficiency, not bias. Moreover, even for variance minimization, the working covariance structure $\bm{M}$ does not need to be specified exactly, only up to a constant factor. The features of $\bm{M}$ that are necessary for the selection of the weights are the correlation within (and, if applicable, between) units across periods and the relative variances of the units and, if applicable, time periods. The working covariance matrix can be written as follows:
@@
where $\bm{M}_{r}$ is the working correlation matrix and $\bm{m}_{v}$ is the vector of (relative) variances of the individual observations. If the units are independent, then $\bm{M}_{r}$—and thus $\bm{M}$—will be block-diagonal, with non-zero entries only for the within-unit covariances. Several common structures could then be specified for the within-unit covariances:
  • Independence: assuming independence of observations would require $\bm{M}_{r}$ to be the identity matrix. This would generate a diagonal $\bm{M}$, where the diagonal entries indicate the relative variances of the different unit-period observations.
  • Exchangeable/compound symmetric: assuming compound symmetry in the within-unit correlation would require $\bm{M}_{r}$ to be block-diagonal with the non-zero off-diagonal entries all equal to the intra-unit (intra-cluster) correlation coefficient, $\rho$.
@@
If the units are exchangeable (e.g., clusters prior to randomization in a SWT), then all of the blocks will be equivalent to one another. More complex variance structures can be implied by specific outcome models as well (see, e.g., Hooper et al. (2016) and Kasza, Hemming, Hooper, Matthews & Forbes (2019)).
@@ 2.7 Estimation and inference
Once the weight vector, $\bm{w}^{*}$, is determined, the estimator is given by:
$\hat{\theta}^{*}=\bm{w}^{*T}\bm{d}=\bm{w}^{*T}\bm{A}\bm{y}.$ (10)
@@
$Var(\hat{\theta}^{*})=\bm{w}^{*T}\bm{A}\widehat{Var}(\bm{y})\bm{A}^{T}\bm{w}^{*},$ (11)
with a sampling distribution depending on the assumed or estimated distribution of $\bm{y}$. This would likely arise from an asymptotic sampling- or model-based approach to inference. This approach has been predominant in the observational staggered adoption literature, with recent work identifying more reasonable assumptions and approaches suitable for dependent units and smaller numbers of units; see Roth et al. (2023) for a summary. Model-based approaches to inference have also been common in the randomized stepped-wedge literature, usually as some form of the linear mixed effects model (see, e.g., Hussey & Hughes (2007) and Hooper et al. (2016)), with robust extensions and misspecification discussed by, e.g., Hughes et al. (2020), Voldal et al. (2022), and Wang, Wang & Li (2024).
Alternatively, inference can proceed non-parametrically using permutation inference. This approach has its basis both in randomization tests used in the SWT case as in Wang & De Gruttola (2017), Thompson et al. (2018), and Kennedy‐Shaffer et al. (2020), and in placebo tests in the staggered adoption case as in Hagemann (2019), MacKinnon & Webb (2020), Shaikh & Toulis (2021), and Roth & Sant’Anna (2023)). This design-based approach to inference is used to test a sharp null hypothesis and conditions on the observed data, using the randomization or “as-if” randomization of the order as the basis for hypothesis testing. Applying the same estimator to the data set under a large number of permutations of the order of adoption provides a null distribution for the estimator to which the observed estimate can be compared. In the remainder of this paper, I focus on this design-based permutation inference approach.
@@ 2.8 Algorithm for construction of estimator
In summary, estimation and inference proceed by the following process.
@@
    Based on the number of periods $J$ and the number of units $N$, identify the matrix $\bm{A}$ that converts the observations $\bm{y}$ to the two-by-two DID estimators $\bm{d}$.
@@
    Find the space of weight vectors $\bm{w}$ such that $\bm{F}^{T}\bm{w}=\bm{v}$ and, if relevant, restrict to the set of weights that give unique estimators.
@@
    Determine an appropriate working covariance matrix $\bm{M}$ (up to a scalar) using subject-matter knowledge, pilot study data, or an assumed outcome model.
@@ 3 Toy Example
I consider several different assumption settings. Throughout, the DID building blocks ensure that weights on rows and columns sum to 0, as discussed by Matthews & Forbes (2017).
3.1 Assumption S5: Homogeneity
@@
  • If $x+y=1$, then the estimator is simply the type-2 estimator $D_{1,2,1,2}$, the “clean” comparison as described by Goodman-Bacon (2021) or the “crossover” estimator CO-1 proposed by Kennedy‐Shaffer et al. (2020);
@@
Note that the purely vertical estimators described by Matthews & Forbes (2017) and Thompson et al. (2018) cannot be constructed as a weighted average of the two-by-two DID estimators.
Assuming independent and homoskedastic observations, $\bm{M}=\bm{I}_{6}$, the six-by-six identity matrix. The constrained minimization of the variance, then, gives the solution $x+y=1/2$ and the estimator $\hat{\theta}^{*}=\frac{D_{1,2,1,2}-D_{1,2,2,3}}{2}$. As mentioned, this corresponds to the “centrosymmetrized” estimator discussed by Matthews & Forbes (2017) and Bowden et al. (2021); the latter proves that this estimator has minimum variance under these assumptions. The same estimator minimizes the variance under a working correlation assumption that is compound symmetric or AR(1), regardless of the correlation value, as long as units are exchangeable and independent.
3.2 Assumption S4: Calendar-Time Heterogeneity
@@
3.3 Assumption S3: Exposure-Time Heterogeneity
@@
4 Comparisons to Existing Estimators
Several existing estimators can be considered under this framework as well, as either weighted averages of certain two-by-two comparisons or, more generally, weighted averages of observations. Goodman-Bacon (2021) discusses this in general for the two-way fixed effects model in the non-randomized staggered adoption setting, while Matthews & Forbes (2017) and Lindner & McConnell (2021) discuss it for the stepped-wedge setting. Here, I consider several alternative estimators that have been proposed for this setting and their relationship to the generalized estimator. The first group were designed for the staggered adoption setting- (see Arkhangelsky & Imbens (2024) for a summary), the second group specifically target the effect in the first period on treatment, and the last estimator uses only vertical comparisons. Comparisons can also be made to methods designed to assess heterogeneity using DID approaches as in Shahn & Hatfield (2024) and Xu et al. (2024).
4.1 Staggered Adoption Methods
Until recently, staggered adoption panel data analyses often used the two-way fixed-effects model (TW). Goodman-Bacon (2021) demonstrated that this could be expressed as a weighted average of two-by-two DID comparisons. Considering each pair of units and periods as a unique comparison (as opposed to the Goodman-Bacon decomposition itself, which combines across timing groups), this method equally weights all comparisons where one unit switches and the other does not, with negative weights for those where the earlier-switching unit is always-treated and the later-switching unit switches. This implicitly makes assumption S5, full treatment effect homogeneity. The TW method is equivalent to the generalized DID with assumption S5 under the assumption that observations are independent and homoskedastic, giving a highly efficient, symmetrized estimator.
Callaway & Sant’Anna (2021) propose a method using only “clean” two-by-two comparisons (CS); i.e., restricting to comparisons where unit $i^{\prime}$ is untreated in both periods $j$ and $j^{\prime}$ (Types 1–3 in Table 1). This can be further restricted to always comparing to a set of never-treated or last-treated control units, but we consider the more general version here. Moreover, they only include comparisons where the baseline period is the last untreated period of unit $i$; i.e., two-by-two components $D_{i,i^{\prime},T_{i}-1,j^{\prime}}$ where $T_{i}\leq j^{\prime}<T_{i^{\prime}}$. These are combined by equal weights into group-time ATTs (where group is defined as the set of units with the same treatment adoption time), which can then be weighted to the desired estimand. Four proposed summary estimates are the “simple” ATT (averaging across group-times proportional to the number of units in each group), the “dynamic” ATT (averaging by time-on-treatment and then weighting equally across those exposure times), the “group” ATT (averaging within groups, and then equally across groups), and the “calendar” ATT (averaging within calendar time periods, and then equally across periods). These all imply different weights on the included two-by-two DID estimators.
Sun & Abraham (2021) similarly propose a method using only “clean” comparisons (SA), specifically comparisons against the never-treated or last-treated group where the last untreated period of unit $i$ is used as the baseline; i.e., two-by-two components $D_{i,i^{\prime},T_{i}-1,j^{\prime}}$, where $j^{\prime}>T_{i}-1$ and $i^{\prime}\in C$, the set of units with the latest treatment adoption time ($T_{i^{\prime}}=\max_{i}T_{i}$ for $i^{\prime}\in C$). These are combined, weighting equally across units within each time period $j^{\prime}$, and then equally weighted across periods. Further restricting the comparison set limits the two-by-two DID estimators that get non-zero weights.
The above methods target a specific estimand by both limiting the comparisons to only allow comparisons with an expectation equal to the desired group-time treatment effect (allowing heterogeneity by both calendar time and exposure time) and then by weighting these together in particular ways. Both can be expressed as weighted averages of two-by-two DID components, as in the generalized DID framework proposed here. Identifying the same target effects can be done more efficiently, however, by incorporating additional components that do not change the expectation but can reduce variance. The tradeoff for this more general approach is that a stronger parallel trends assumption may be required, as more cluster-periods are used in the estimation.
4.2 First-Period Effect Methods
The method proposed by de Chaisemartin & D’Haultfœuille (2020) (CH) allows comparisons to not-yet-treated groups, but only one-period ahead differences; i.e. only Type 2 comparisons $D_{i,i^{\prime},T_{i}-1,T_{i}}$, where $T_{i^{\prime}}>T_{i}$. These are then averaged within each timing group, and then across groups, weighted equally. This inherently restricts the target estimand to the effect of treatment in the first period of adoption.
Similarly, the crossover estimator proposed by Kennedy‐Shaffer et al. (2020) only considers first-period effects. CO-1 weights equally across timing groups, and is thus equal to the de Chaisemartin & D’Haultfœuille (2020) estimator. CO-2 uses the same comparisons, but different weights, weighting each timing group proportional to the harmonic mean of the number of treated and control units. CO-3 is similar, but allows always-treated units to act as controls as well, incorporating Type 5 comparisons as well. Weighting is similar to CO-1. Again, these all target the first-period effect, but with different weights having different properties under calendar-time or unit heterogeneity (CO-3 in particular requires no dynamic effects).
These two methods target a very specific estimand, the first-period effect of adoption, and are best suited to ignoring or averaging over calendar time heterogeneity. CO-3 has improved efficiency under the assumption of no dynamic treatment effects because it uses more data (see, e.g., Bowden et al. (2021)), but can become biased for the first-period effect if that assumption does not hold (Kennedy‐Shaffer et al. (2020)). The generalized DID method also can be designed to target this effect alone; it will generally use more comparisons and be more efficient under any specific assumption setting. However, it loses the simplicity of restricting the set to just include those comparisons that clearly target that specific effect which in non-randomized settings may require a stronger parallel trends assumption.
4.3 Observation Weight Method
Finally, the non-parametric within-period method (NP) proposed by Thompson et al. (2018) uses individual observations as the components instead of two-by-two DID comparisons. It compares treated to untreated units within each calendar period, and then averages these comparisons across periods, with weights either equal across periods, proportional to the number of treated groups, or inversely proportional to the estimated variance of the comparison. For the inverse-variance weights, we assume equal variance across observations rather than using the sample variance, so that weights can be computed without the observation data. The within-period approach is particularly useful when calendar-time heterogeneity is expected to be more pronounced or of greater interest than exposure-time heterogeneity. However, the lack of horizontal comparisons limits the power and efficiency of the method (see, e.g., Matthews & Forbes (2017), Thompson et al. (2018), and Kennedy‐Shaffer et al. (2020)).
4.4 Comparisons
Observation weights and, for all except the final method described, DID weights for these methods are implemented in the Comp_Ests_Weights function in the R code, to enable investigators to compare the weights used in these different approaches to those under the generalized DID framework. These weights can all be calculated a priori, so that the advantages and disadvantages of the weighting schemes can be considered without being influenced by the observed data. Overall, the generalized DID method tends to use more information to target similar effects and allows more flexibility for the researcher. This may result in reduced variance and improved efficiency of estimation, especially under more effect homogeneity and if the working covariance matrix is correctly specified. The tradeoff, however, is that it is generally a more complex estimator than the ones proposed here, with less clearly interpretable weights. By using more comparisons, it also requires the assumptions made (e.g., parallel trends and any homogeneity assumptions) to hold true for a broader set of units and periods.
@@ 5 Data Example: Stepped Wedge Trial
This study randomized fourteen clusters into seven treatment sequences, with observations in each of eight months. The first sequence, consisting of two clusters, began the intervention in the second month; each subsequent month, two more clusters began the intervention (Durovni et al. (2014)). The patient outcomes, namely unsuccessful treatment of the patient defined as any outcome other than “cure and treatment completion without evidence of failure” (Trajman et al. (2015)), were retrieved using the Brazilian national tuberculosis information system over one year after diagnosis.
The original analysis found a reduction in this composite outcome, although it was not statistically significant: an odds ratio of 0.93 with a 95% confidence interval of 0.79–1.08 (Trajman et al. (2015)). Accounting for time trends in the analyses, however, re-analysis by Thompson et al. (2018) using the purely vertical non-parametric within-period method found a larger and statistically significant effect: an odds ratio of 0.78 with a p-value of 0.02. Similar results were found by Kennedy‐Shaffer et al. (2020) using the crossover method, which estimated an odds ratio of 0.70 with a permutation-based p-value of 0.046. These results indicate the possible impacts of different treatment effect heterogeneity assumptions and different choices of estimand.
I re-analyze these data using several estimators from the generalized DID framework, beginning with four overall treatment effects:
@@
  • $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}^{(4)}_{j}$, the average period-specific treatment effect across months 2–7 of the trial under calendar-time heterogeneity (assumption S4);
  • $\frac{1}{7}\sum_{a=1}^{7}\hat{\theta}^{(3)}_{a}$, the average time-on-treatment effect across exposure times 1–7 under exposure-time heterogeneity (assumption S3); and
  • $\frac{1}{21}\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{j,a}$, the average across all identifiable calendar- and exposure-time combinations of the trial under both heterogeneities (assumption S2).
Note that the period-specific treatment effects in month 8 (the last month of the study) are not identifiable since all clusters were in the treated condition at that point. These clusters do contribute to estimation of $\hat{\theta}^{(5)}$ and $\hat{\theta}^{(3)}_{7}$, however, under assumptions S5 and S3, as those assumptions allow us to borrow information across different calendar times in estimating common treatment effects. The odds ratio (calculated using log-odds of events in each cluster-period as the outcome) and risk difference (calculated using proportion of events in each cluster-period as the outcome) estimates and permutation-test p-values (using 1000 permutations) are shown in Table 4. All results shown were calculated using an exchangeable (within-cluster) variance structure with intracluster correlation coefficient $\rho=0.003$ as estimated by Thompson et al. (2018). Results under independent and auto-regressive variance structures are quite similar and shown in Appendix D.
@@
S5 $\hat{\theta}^{(5)}$ 0.771 0.013 -7.35% 0.020
S4 $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}^{(4)}_{j}$ 0.784 0.021 -6.81% 0.054
S3 $\frac{1}{7}\sum_{a=1}^{7}\hat{\theta}^{(3)}_{a}$ 0.828 0.413 -5.57% 0.456
S2 $\frac{1}{21}\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{j,a}$ 0.801 0.149 -6.11% 0.229
@@
To understand the form of the estimator, we can also examine the weights it gives to each observation (note: the weights on each two-by-two DID estimator are also available, but less interpretable as many different weight vectors can yield the same weights on the observations). The weights for the estimator $\hat{\theta}^{(5)}$ are shown in the heat map in Figure 1. Pairs of clusters (e.g., 1 and 2) have nearly identical weights because they are in the same sequence and assumed to be exchangeable. These weights can be compared to the information content of specific cells under different heterogeneity and correlation assumptions as proposed by Kasza & Forbes (2019) and Kasza, Taljaard & Forbes (2019). Note heuristically that in this case, the largest (in magnitude) weights are near the diagonal, followed by the top-right and bottom-left cells, which are also the highest information-content cells as shown in Figure 4 of Kasza & Forbes (2019). In addition, the weights satisfy the symmetry discussed by Matthews & Forbes (2017) and Bowden et al. (2021), which maximizes efficiency. Slight differences arise when assumption S4 is used instead with the calendar time-averaged estimator; Figure 5 displays these weights.
@@
This method provides flexibility to investigate other treatment effects in the heterogeneous settings as well. Considering calendar-time effects under assumption S4, for example, one can estimate the treatment effect in each of months 2 through 7, $\hat{\theta}^{(4)}_{2},\ldots,\hat{\theta}^{(4)}_{7}$. For additional robustness against exposure-time heterogeneity, one can also estimate month-specific treatment effects under assumption S2 by averaging across exposure times within each of those months: $\hat{\theta}^{(2)}_{j}\equiv\frac{1}{j-1}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{a,j}$ for all $j=2,\ldots,7$. The estimation and permutation test results on both the multiplicative (odds ratio) and additive (risk difference) scales are shown in Table 5. These results can be compared to the results in Table 2 of Thompson et al. (2018), which reports period-specific risk differences estimated using the purely vertical non-parametric within-period estimator (note that those are labelled by treatment periods, one less than the corresponding month numbers given here). Differences between the methods arise as the results presented here use horizontal information as well, but target the same estimand. Note that no p-values are adjusted for multiple testing.
Table 5: Odds ratio and risk difference estimates and permutation test p-values for generalized DID estimators of calendar-time treatment effects of XpertMTB/RIF testing on tuberculosis outcomes in Brazil, 2012, using an exchangeable variance structure with ICC 0.003.